OpenStudy (anonymous):

\[x^2 +\sqrt{xy}=xy^5\] compute y' pls show all steps involved

OpenStudy (anonymous):

steps are as if \[y=f(x), y'=f'(x)\] , chain rule, product rule, etc \[x^2 +\sqrt{xy}=xy^5\] \[2x+\frac{xy'+y}{2\sqrt{xy}}=y^5+5xy^4u'\] then algebra to solve for y'

OpenStudy (anonymous):